OpenStudy (anonymous):

x^2-28x What number must you add to the polynomial below to complete the square? Please show steps ! Please ! Thanks

OpenStudy (anonymous):

lets say you add b to make it some (x-a)^2 then x^2 - 28x +b = (x-a)^2 x^2 -28x + b= x^2 -2ax + a^2 Now compare coefficients -28=-2a and b=a^2 so a=14 b=196

OpenStudy (anonymous):

So which number would I add to complete the square ? Also I think your solving for a and b . i just need to know what number makes this equation a perfect square binomial.

OpenStudy (anonymous):

didnt i mention that you add b?

OpenStudy (anonymous):

196 right ?

OpenStudy (anonymous):

right